Problem
A boxing ring is in the shape of a square 20ft on each side. How far apart are the firefighters when thay are in opposite corners of the ring?

Thank

Answer by stanbon(75887)

A boxing ring is in the shape of a square 20ft on each side. How far apart are the firefighters when thay are in opposite corners of the ring?
----------------
Use Pythgoras:
Let the distance between opposite corners be "x".
x^2 = 20^2 + 20^2
x^2 = 2*20^2
Take the square root to get:
x = 20*sqrt(2)
x is approximately 28.28 ft.
=======================
Cheers,
Stan H.